OpenStudy (anonymous):

use the foil method to find the product. (7x^5-4)(x^8-2)

OpenStudy (amistre64):

what is the foil method?

OpenStudy (anonymous):

i dont remember

OpenStudy (anonymous):

first terms outter terms inner terms last terms multiply those together then combine like terms and you have your answer.

OpenStudy (amistre64):

its rather ironic that foil is a memory device that you cant remember :)

OpenStudy (amistre64):

gt was right :)

OpenStudy (anonymous):

FOIL is a ridiculous memory based method to teach "distributive" property of factors.

OpenStudy (anonymous):

(a+b)(c+d) = ac + ad + bc + bd according to FOIL.

OpenStudy (anonymous):

i know when im doing a problem like that im always saying first outer inner last while im doing the problem lol

OpenStudy (anonymous):

FOIL focuses on the "order". It is mathematically WRONG to focus on the order of the terms. Order does not matter - mathematically speaking.

OpenStudy (amistre64):

i just multiply them together like they taught me back in teh 2nd grade or so

OpenStudy (anonymous):

(7x^5-4)(x^8-2) = 7x^5 * x^8 - 2*7*x^5 - 4*x^8 + 4*2 = 7x^13 - 14x^5 - 4x^8 + 8

OpenStudy (anonymous):

thats what i got to thank you

OpenStudy (anonymous):

what answer did you get?

OpenStudy (anonymous):

Another thing that annoys me a lot is the "soh cah toa" in trigonometry. It drives me bananas.

OpenStudy (anonymous):

i have yet done trigonometry so i have no clue what that is lol

OpenStudy (anonymous):

i have not yet i mean

OpenStudy (anonymous):

its algibra

OpenStudy (anonymous):

algebra?

OpenStudy (anonymous):

Thank god you don't. If you did, you would be annoyed too..... :)

OpenStudy (anonymous):

lol i stayed away from it in high school cause it looked way to hard

OpenStudy (anonymous):

It is EASY. Simply ratio of sides of triangles = trigonometry.

OpenStudy (anonymous):

sounds like a indian name lol

OpenStudy (anonymous):

Greek I think. :) All same....complex names.

OpenStudy (anonymous):

ill have to google it just to see how to do it

OpenStudy (anonymous):

ok well im off to do my work. i just stopped in to see who i could help. and i helped ohhh 2 people lol have a good night

OpenStudy (anonymous):

Cool. Good luck and good night.

OpenStudy (anonymous):

thanks :)